# Generalized Out-of-Distribution Detection and Beyond in Vision Language Model Era: A Survey

### v1

@@ 4.1 Zero-shot Out-of-Distribution Detection
Zero-shot OOD detection was proposed in 2021 by Fort et al. [38]. Since then, a growing number of methods have been proposed year by year.
Definition of Zero-shot OOD Detection  In zero-shot OOD detection, the term “Zero-shot” refers to the non-use of ID images during both training and inference phases. For instance, the method with additional training with auxiliary datasets (non-use of ID images) can be regarded as a zero-shot method [40]. The method with the pre-processing of the ID class texts can also be regarded as a zero-shot method [38, 39, 45, 48].
@@ 4.1.1 Training-free Methods
a. With OOD Prompts  CLIP-based OOD detection started in this setting. The earliest work is ZeroOE [38]. ZeroOE feeds the potential OOD labels to the textual encoder of CLIP. However, the method of using known OOD labels is infeasible for real-world applications. To solve this problem, ZOC [39] proposed to train an OOD label generator based on the visual encoder of CLIP and use the generated pseudo-OOD labels for OOD detection. However, when dealing with large-scale datasets encompassing a multitude of ID classes, the label generator may not generate effective candidate OOD labels, resulting in poor performance. Building on these early works [38, 39], recent works focus on how to obtain high-quality OOD labels through either (i) OOD label retrieval [45, 100] or (2) OOD label generation [48]. (i) One of the representative retrieval-based methods is NegLabel [45]. NegLabel selects high-quality OOD labels from extensive corpus databases by calculating the distance between an extracted OOD label and ID label. (ii) One of the representative generation-based methods is EOE [48]. EOE utilizes Large Language Models (LLMs) to produce high-quality OOD labels. By modifying the prompts given to the LLM, EOE can be generalized to a variety of tasks, including far and near OOD detection.
@@
b. Without OOD Prompts  In zero-shot OOD detection, many methods utilize OOD labels, but the difficulty and cost of creating these labels pose challenges. To address these issues, Ming et al. [26] proposed MCM, which uses only ID labels to detect OOD. MCM is a simple approach that devises softmax scaling to align visual features with textual concepts for OOD detection. Despite its simplicity, MCM has high effectiveness and scalability, and it serves as a crucial baseline in CLIP-based OOD detection. Building on the concept of MCM, Miyai et al. [96] proposed GL-MCM, which extends MCM by just adding a local MCM score to enhance the fine-grained detection capability in local regions. SeTAR [97] enhances MCM and GL-MCM by changing the model’s weight matrices using a simple greedy search algorithm. We consider these methods to be post-hoc methods for CLIP-based OOD detection in that they directly employ an ID classifier for OOD detection. Due to their simplicity and high scalability, these post-hoc methods can bring fundamental performance improvements for many subsequent methods [28, 44, 47]. Therefore, we expect that this field should be developed further in the future, reflecting the trajectory of the field before CLIP emerged [101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111].
@@ 4.1.2 Auxiliary Training-based Methods
CLIPN [40] is the only auxiliary training-based method for zero-shot OOD detection. CLIPN aims to empower the logic of saying “no” within CLIP, and it designs a novel learnable “no” prompt and an additional “no” text encoder to capture negation semantics within images. To create an additional text encoder, CLIPN needs to be pre-trained on the CC-3M dataset [112]. While the extensive pre-training of CLIPN may indeed lead to intensive computations and lower scalability, once pre-trained, it performs zero-shot OOD detection across a wide range of domains with comparable performance to few-shot OOD detection methods [28, 44]. In the future, within this field, there is potential for zero-shot open-vocabulary OOD detection to further advance the realm of zero-shot OOD detection, which will be discussed later in Sec. 7.2.
@@ 4.2 Few-shot Out-of-Distribution Detection
Few-shot OOD detection was concurrently proposed by Miyai et al. [28] and Ming et al. [41] in June 2023. Since then, it has become the most active research area in CLIP-based OOD detection.
Definition of Few-shot OOD Detection  CLIP-based few-shot OOD detection aims to detect OOD images using only a few labeled ID images. In few-shot OOD detection, the term “Few-shot” refers to the use of a few ID images during training or inference phases. For instance, the method with additional training with a few ID images can be regarded as a few-shot method [33]. Even without training, if a method uses a few ID images as a reference, we regard it as a few-shot method [27]. Regarding the number of shots, it is common to experiment with 1-shot to 16-shot [28, 99], following the closed-set setting [24].
@@ 4.2.1 ID Training-based Methods
a. Without OOD Prompts  Few-shot OOD detection began in this setting. Ming et al. [41] proposed PEFT-MCM for CLIP-based OOD detection, which demonstrates the effectiveness of combining parameter-efficient tuning methods (e.g., prompt learning [24] or adapter [113]) and MCM [26]. Concurrently, Miyai et al. [28] proposed LoCoOp, a pioneer prompt learning approach for few-shot OOD detection. LoCoOp enhances CoOp’s [24] OOD detection capabilities by performing OOD regularization with local OOD features. LoCoOp is the simplest prompt learning method and serves as a crucial baseline in few-shot OOD detection. Unlike LoCoOp, which utilizes non-ID local regions for OOD regularization, GalLoP [98] proposes an approach that utilizes local ID regions to enable a more fine-grained distinction between ID and OOD samples. GalLoP learns a diverse set of prompts by utilizing both global and local visual representations, thereby enhancing the detection capabilities.
@@
b. With OOD Prompts  Similar to zero-shot OOD detection, recent works in few-shot OOD detection utilize additional OOD prompts [44, 46, 47]. As representative methods, LSN [44] and NegPrompt [47] were proposed concurrently. They state that the simple negative prompts added “not” (e.g., “not a photo of a [cls]”) fail to capture the dissimilarity for identifying OOD samples. Therefore, by preparing negative prompts and training with them, LSN and NegPrompt can learn suitable negative prompts, enabling more accurate detection of OOD samples. The difference between LSN and NegPrompt lies in their approach to the use of negative prompts. LSN prepares unique negative prompts for each class and learns suitable negative prompts for each class. In contrast, NegPrompt prepares multiple negative prompts common to all ID classes and trains them to learn generic templates representing the negative semantics of any given class labels. Also, NegPrompt tested the performances in the hard OOD detection setting with ImageNet-protocol [54], outperforming LoCoOp and CoOp. Alternatively, IDPrompt [46] takes a different approach by introducing ID-like prompts, which are designed for OOD features that are close to the ID features. It extracts ID-like OOD regions in ID training images and trains ID-like prompts with these extracted OOD data. In a unique direction, LAPT [114] proposes an automatic sample collection strategy that retrieves or generates training ID images only with ID class names, which achieves high performance without image collection and annotation costs. LAPT then performs distribution-aware prompt learning, which distinguishes between ID class and OOD class tokens. LAPT is positioned within the context of more efficient few-shot OOD detection in this survey paper since it requires generating or retrieving “ID images” for the data collection.
@@
In the context of few-shot OOD detection, recently, Li et al. [47] proposed a new problem setting called open-vocabulary OOD (OV-OOD) detection. While common few-shot OOD detection involves training on images from all ID classes during training, OV-OOD detection involves training on images from just a small subset of ID classes and performing OOD detection using all ID classes at inference time. Formally, we define a subset of semantic labels $\mathcal{Y}_{\text{ID,sub}}\subset\mathcal{Y}_{\text{ID}}$, where $\mathcal{Y}_{\text{ID}}$ represents all ID labels. Based on this subset of labels, we define a corresponding subset dataset $\mathcal{D}_{\text{ID,sub}}^{\texttt{train}}\subset\mathcal{D}_{\text{ID}}^{\texttt{train}}$. During training, only $\mathcal{D}_{\text{ID,sub}}^{\texttt{train}}$ is used. Then, at inference time, all ID classes $\mathcal{Y}_{\text{ID}}$ are used, and the goal is to detect OOD from a combination of all ID test data $\mathcal{D}_{\text{ID}}^{\texttt{test}}$ with $\mathcal{Y}_{\text{ID}}$ and OOD test data $\mathcal{D}_{\text{OOD}}^{\texttt{test}}$ with $\mathcal{Y}_{\text{OOD}}$. For this setting, existing few-shot OOD detection methods [28, 47] can be easily applied by simply combining the rest of the ID classes. In particular, NegPrompt [47] learns general negative prompts that are not specific to the training ID classes, so it achieves high performance in OV-OOD detection.
@@ 4.2.2 Training-free Methods
Training-free few-shot OOD detection is a novel research field, and only Dual-Adapter [99] falls under this category. Dual-Adapter adopts a prior-based method Tip-Adapter [113], which leverages both textual and visual features with a cache model and enhances performance without training. To adapt this to few-shot OOD detection, Dual-Adapter employs the concept of dual cache modeling and constructs Positive-Adapter and Negative-Adapter, and identifies OOD samples with the prediction difference with both adapters.
4.3 Other Important Research Directions
4.3.1 CLIP-based Full-spectrum OOD Detection
@@
CLIP-based full-spectrum OOD (FS-OOD) detection is a crucial challenge [115]. FS-OOD detection was proposed by Yang et al. [116] in 2022 as an important setting that considers both OOD generalization [117, 118] and OOD detection simultaneously. Unlike standard OOD detection, which only focuses on semantic shifts between training and test distributions, FS-OOD detection further considers non-semantic covariate shift by including covariate-shifted ID images. As for the benchmarks, OpenOOD v1.5 [119] provides two large-scale benchmarks based on ImageNet-200 and ImageNet-1K, incorporating ImageNet-C [117] with image corruptions, ImageNet-R [118] with style changes, and ImageNet-V2 [120] with resampling bias as ID. As for CLIP-based methods, LSA [115] uses a bidirectional prompt customization mechanism, which adjusts discriminative ID and OOD boundary.
4.3.2 Other Tasks with CLIP-based OOD Detection
@@
Unsupervised Universal Fine-Tuning  CLIP-based OOD detection is useful for a new task called Unsupervised Universal Fine-Tuning (UUFT) [49]. UUFT is a problem of unsupervised learning for outlier detection (OD). Existing studies for unsupervised learning assumed that all unlabeled images belong to one of the ID classes [121, 122, 123], but they require prior knowledge of exact class names linked to ground truth labels, which restricts their usefulness in various real-world situations. For a more realistic setting, UUFT assumes that OOD images are included in the unlabeled images. To detect OOD images during training, they developed MCM [26] and proposed UEO, which leverages sample-level confidence to approximately minimize the conditional entropy of confident instances and maximize the marginal entropy of less confident instances.
@@
Open-world Prompt Tuning  CLIP-based OOD detection is useful for a new task called Open-world Prompt Tuning [124]. Open-world Prompt Tuning is a task that evaluates the classification accuracy on a mix of known and novel ID classes while training the model with known classes. To solve this problem, Zhou et al. [124] proposed DeCoOp which incorporates OOD detection into the inference pipelines and improves the base-to-new separability, preventing performance degradation on new classes.
5 CLIP-based AD: Methodology

### v2

@@ 4.1 Zero-shot Out-of-Distribution Detection
Zero-shot OOD detection was proposed in 2021 by Fort et al. (2021). Since then, a growing number of methods have been proposed year by year.
Definition of Zero-shot OOD Detection  In zero-shot OOD detection, the term “Zero-shot” refers to the non-use of ID images during both training and inference phases. For instance, the method with additional training with auxiliary datasets (non-use of ID images) can be regarded as a zero-shot method (Wang et al., 2023a). The method with the pre-processing of the ID class texts can also be regarded as a zero-shot method (Fort et al., 2021; Esmaeilpour et al., 2022; Jiang et al., 2024; Cao et al., 2024a).
@@ 4.1.1 Training-free Methods
a. With OOD Prompts  VLM-based OOD detection started in this setting. The earliest work is ZeroOE (Fort et al., 2021). ZeroOE feeds the potential OOD labels to the textual encoder of CLIP. However, the method of using known OOD labels is infeasible for real-world applications. To solve this problem, ZOC (Esmaeilpour et al., 2022) proposed to train an OOD label generator based on the visual encoder of CLIP and use the generated pseudo-OOD labels for OOD detection. However, when dealing with large-scale datasets encompassing a multitude of ID classes, the label generator may not generate effective candidate OOD labels, resulting in poor performance. Building on these early works (Fort et al., 2021; Esmaeilpour et al., 2022), recent works focus on how to obtain high-quality OOD labels through either (i) OOD label retrieval (Jiang et al., 2024; Ding & Pang, 2024) or (2) OOD label generation (Cao et al., 2024a). (i) A representative retrieval-based method is NegLabel (Jiang et al., 2024). NegLabel selects high-quality OOD labels from extensive corpus databases by calculating the distance between an extracted OOD label and ID label. AdaNeg (Zhang & Zhang, 2024) extends NegLabel by incorporating adaptive negative proxies, which are dynamically generated during inference by leveraging actual OOD images. This adaptation ensures a closer alignment with the test-time OOD label space, thereby enhancing the effectiveness of negative proxy guidance. Chen et al. (2024a) proposed Conjugated Semantic Pool (CSP), which enhances OOD detection by expanding the semantic space with modified superclass names. (ii) A representative generation-based method is EOE (Cao et al., 2024a). EOE utilizes Large Language Models (LLMs) to produce high-quality OOD labels. By modifying the prompts given to the LLM, EOE can be generalized to a variety of tasks, including far and near OOD detection.
@@
b. Without OOD Prompts  In zero-shot OOD detection, many methods utilize OOD labels, but the difficulty and cost of creating these labels pose challenges. To address these issues, Ming et al. (2022a) proposed MCM, which uses only ID labels to detect OOD. MCM is a simple approach that devises softmax scaling to align visual features with textual concepts for OOD detection. Despite its simplicity, MCM has high effectiveness and scalability, and it serves as a crucial baseline in VLM-based OOD detection. Building on the concept of MCM, Miyai et al. (2025b) proposed GL-MCM, which extends MCM by just adding a local MCM score to enhance the fine-grained detection capability in local regions. Based on MCM and GL-MCM SeTAR (Li et al., 2024e) enhances them by changing the model’s weight matrices using a simple greedy search algorithm. As for another direction, some works adopt data-centric approaches (Liu & Christopher, 2024; Zhang et al., 2024a).  Liu & Christopher (2024) proposed TAG, which applies simple augmentations to the ID text prompts.  Zhang et al. (2024a) proposed CoVer, which enhances separability by averaging confidence scores across corrupted inputs, leveraging the greater confidence drop in OOD samples under corruptions.
We consider these methods to be post-hoc methods for VLM-based OOD detection in that they directly employ an ID classifier for OOD detection. Due to their simplicity and high scalability, these post-hoc methods can bring fundamental performance improvements for many subsequent methods (Miyai et al., 2023b; Nie et al., 2023; Li et al., 2024b). Therefore, we expect that this field should be developed further in the future, reflecting the trajectory of the field before CLIP emerged (Hendrycks & Gimpel, 2017; Liang et al., 2018; Lee et al., 2018; Liu et al., 2020; Sastry & Oore, 2020; Wang et al., 2021; Zhang et al., 2023; Sun & Li, 2022; Sun et al., 2022; Lin et al., 2021; Sastry & Oore, 2019).
@@ 4.1.2 Auxiliary Training-based Methods
CLIPN (Wang et al., 2023a) is the only auxiliary training-based method for zero-shot OOD detection. CLIPN aims to empower the logic of saying “no” within CLIP, and it designs a novel learnable “no” prompt and an additional “no” text encoder to capture negation semantics within images. To create an additional text encoder, CLIPN needs to be pre-trained on the CC-3M dataset (Sharma et al., 2018). While the extensive pre-training of CLIPN may indeed lead to intensive computations and lower scalability, once pre-trained, it performs zero-shot OOD detection across a wide range of domains with comparable performance to few-shot OOD detection methods (Miyai et al., 2023b; Nie et al., 2023).
@@ 4.2 Few-shot Out-of-Distribution Detection
Few-shot OOD detection was concurrently proposed by Miyai et al. (2023b) and Ming & Li (2024a) in June 2023. Since then, it has become an active research area in VLM-based OOD detection.
Definition of Few-shot OOD Detection  VLM-based few-shot OOD detection aims to detect OOD images using only a few labeled ID images. In few-shot OOD detection, the term “Few-shot” refers to the use of a few ID images during training or inference phases. For instance, the method with additional training with a few ID images can be regarded as a few-shot method (Li et al., 2024c). Even without training, if a method uses a few ID images as a reference, it is treated as a few-shot method (Jeong et al., 2023). Regarding the number of shots, it is common to experiment with 1-shot to 16-shot (Miyai et al., 2023b; Chen et al., 2024b), following the closed-set setting (Zhou et al., 2022c).
@@ 4.2.1 ID Training-based Methods
a. Without OOD Prompts  Few-shot OOD detection began in this setting. Ming & Li (2024a) proposed PEFT-MCM for VLM-based OOD detection, which demonstrates the effectiveness of combining parameter-efficient tuning methods (e.g., prompt learning (Zhou et al., 2022c) or adapter (Zhang et al., 2022)) and MCM (Ming et al., 2022a). Concurrently, Miyai et al. (2023b) proposed LoCoOp, a pioneer prompt learning approach for few-shot OOD detection. LoCoOp enhances CoOp’s (Zhou et al., 2022c) OOD detection capabilities by performing OOD regularization with local OOD features. LoCoOp is the simplest prompt learning method and serves as a crucial baseline in few-shot OOD detection. To improve the performance of LoCoOp, Self-Calibrated Tuning (SCT) (Yu et al., 2024) dynamically adjusts the weight of OOD regularization based on prediction uncertainty to prevent unreliable OOD features from degrading the model’s performance. Unlike LoCoOp, which utilizes non-ID local regions for OOD regularization, GalLoP (Lafon et al., 2024) proposes an approach that utilizes local ID regions to enable a more fine-grained distinction between ID and OOD samples. GalLoP learns a diverse set of prompts by utilizing both global and local visual representations, thereby enhancing detection capabilities.
@@
b. With OOD Prompts  Similar to zero-shot OOD detection, recent works in few-shot OOD detection utilize additional OOD prompts (Nie et al., 2023; Bai et al., 2024a; Li et al., 2024b). As representative methods, LSN (Nie et al., 2023) and NegPrompt (Li et al., 2024b) were proposed concurrently. They state that the simple negative prompts added “not” (e.g., “not a photo of a [cls]”) fail to capture the dissimilarity for identifying OOD samples. Therefore, by preparing negative prompts and training with them, LSN and NegPrompt can learn suitable negative prompts, enabling more accurate detection of OOD samples. The difference between LSN and NegPrompt lies in their approach to the use of negative prompts. LSN prepares unique negative prompts for each class and learns suitable negative prompts for each class. In contrast, NegPrompt prepares multiple negative prompts common to all ID classes and trains them to learn generic templates representing the negative semantics of any given class labels. Additionally, NegPrompt tested the performances in the hard OOD detection setting with ImageNet-protocol (Palechor et al., 2023), outperforming LoCoOp and CoOp. Alternatively, ID-like-Prompt (Bai et al., 2024a) takes a different approach by introducing ID-like prompts, which are designed for capturing OOD features that are close to the ID features. It extracts ID-like OOD regions from ID training images and trains ID-like prompts using these extracted OOD data, enabling the model to capture more subtle differences within images. In a unique direction, LAPT (Zhang et al., 2024d) proposes an automatic sample collection strategy that retrieves or generates training ID images only with ID class names, which achieves high performance without image collection and annotation costs. LAPT then performs distribution-aware prompt learning, which distinguishes between ID class and OOD class tokens. LAPT is positioned within the context of more efficient few-shot OOD detection in this survey paper since it requires generating or retrieving “ID images” for the data collection.
@@
In the context of few-shot OOD detection, recently, Li et al. (2024b) proposed a new problem setting called open-vocabulary OOD (OV-OOD) detection. While common few-shot OOD detection involves training on images from all ID classes during training, OV-OOD detection involves training on images from just a small subset of ID classes and performing OOD detection using all ID classes at inference time. Formally, we define a subset of semantic labels $\mathcal{Y}_{\text{ID,sub}}\subset\mathcal{Y}_{\text{ID}}$, where $\mathcal{Y}_{\text{ID}}$ represents all ID labels. Based on this subset of labels, we define a corresponding subset dataset $\mathcal{D}_{\text{ID,sub}}^{\texttt{train}}\subset\mathcal{D}_{\text{ID}}^{\texttt{train}}$. During training, only $\mathcal{D}_{\text{ID,sub}}^{\texttt{train}}$ is used. Then, at inference time, all ID classes $\mathcal{Y}_{\text{ID}}$ are used, and the goal is to detect OOD from a combination of all ID test data $\mathcal{D}_{\text{ID}}^{\texttt{test}}$ with $\mathcal{Y}_{\text{ID}}$ and OOD test data $\mathcal{D}_{\text{OOD}}^{\texttt{test}}$ with $\mathcal{Y}_{\text{OOD}}$. For this setting, existing few-shot OOD detection methods (Miyai et al., 2023b; Li et al., 2024b) can be easily applied by simply combining the rest of the ID classes. In particular, NegPrompt (Li et al., 2024b) learns general negative prompts that are not specific to the training ID classes, so it achieves high performance in OV-OOD detection.
@@ 4.2.2 Training-free Methods
Recently, training-free few-shot OOD detection methods, which use ID images solely as references during inference, have gained attention. Dual-Adapter (Chen et al., 2024b) adopts a prior-based method Tip-Adapter (Zhang et al., 2022), which leverages both textual and visual features with a cache model and enhances performance without training. To adapt this to few-shot OOD detection, Dual-Adapter employs the concept of dual cache modeling and constructs Positive-Adapter and Negative-Adapter, and identifies OOD samples with the prediction difference with both adapters. Dual-Pattern Matching (DPM) (Zhang et al., 2024f) enhances CLIP for OOD detection by utilizing both textual and visual ID patterns. DPM stores ID class-wise text features as the textual pattern and the aggregated ID visual information as the visual pattern. During inference, it evaluates the similarity of inputs to both patterns to identify OOD samples.
4.3 Other Directions
4.3.1 VLM-based Full-spectrum OOD Detection
@@
VLM-based full-spectrum OOD (FS-OOD) detection is a crucial challenge (Lu et al., 2023). FS-OOD detection was proposed by Yang et al. (2023) as an important setting that considers both the detection of semantic shift and the ability to recognize covariate-shifted data (Hendrycks & Dietterich, 2019; Hendrycks et al., 2021a). Unlike standard OOD detection, which only focuses on semantic shifts between training and test distributions, FS-OOD detection further considers non-semantic covariate shift by including covariate-shifted ID images. As for the benchmarks, OpenOOD v1.5 (Zhang et al., 2024b) provides two large-scale benchmarks based on ImageNet-200 and ImageNet-1K, incorporating ImageNet-C (Hendrycks & Dietterich, 2019) with image corruptions, ImageNet-R (Hendrycks et al., 2021a) with style changes, and ImageNet-V2 (Recht et al., 2019) with resampling bias as ID. As for VLM-based methods, LSA (Lu et al., 2023) uses a bidirectional prompt customization mechanism, which adjusts discriminative ID and OOD boundary. The latest work, Noda et al. (2025) introduces ImageNet-FS-X and Wilds-FS-X as new benchmarks. ImageNet-FS-X utilizes the hierarchical structure of ImageNet labels to define a more challenging ID/OOD split while also evaluating robustness to covariate shifts in ImageNet variants. To further align these benchmarks with real-world testbeds, Wilds-FS-X extends Wilds (Koh et al., 2021) for FS-OOD evaluation. These benchmarks present substantial room for further improvements.
4.3.2 Other Tasks with VLM-based OOD Detection
@@
Unsupervised Universal Fine-Tuning  VLM-based OOD detection is useful for a new task called Unsupervised Universal Fine-Tuning (UUFT) (Liang et al., 2024). UUFT is a problem of unsupervised learning for outlier detection (OD). Existing studies for unsupervised learning assumed that all unlabeled images belong to one of the ID classes (Huang et al., 2022; Shu et al., 2022; Tanwisuth et al., 2023), but they require prior knowledge of exact class names linked to ground truth labels, which restricts their usefulness in various real-world situations. For a more realistic setting, UUFT assumes that OOD images are included in the unlabeled images. To detect OOD images during training, they developed MCM (Ming et al., 2022a) and proposed UEO, which leverages sample-level confidence to approximately minimize the conditional entropy of confident instances and maximize the marginal entropy of less confident instances.
@@
Open-world Prompt Tuning  VLM-based OOD detection is useful for a new task called Open-world Prompt Tuning (Zhou et al., 2024b). Open-world Prompt Tuning is a task that evaluates the classification accuracy on a mix of known and novel ID classes while training the model with known classes. To solve this problem, Zhou et al. (2024b) proposed DeCoOp which incorporates OOD detection into the inference pipelines and improves the base-to-new separability, preventing performance degradation on new classes.
5 VLM-based AD: Methodology